OpenStudy (anonymous):

The average speed of an express train is 15mi/hr faster than the average speed of a local train. It takes the local train 2 hours longer than the express train to cover 360 miles. What is the average speed of the express train?

OpenStudy (anonymous):

speed = dist / time let the local train speed be x x = 360 / 2 + t x + 15 = 360 / t 360/t - 15 = 360 / 2 + t 360 - 15t / t = 360 / 2 + t (t+2) ( 360 - 15t) = 360t -15t^2 - 30t + 720 = 0 t^2 + 2t - 48 = 0 t^2 + 8t -6t -48=0 t(t+8)-6(t+8) = 0 t = 6 or -8 t = 6 x+15 = 360/6 = 60 x = 60 - 15 = 45 miles / hr

OpenStudy (anonymous):

the answer says 60 mi/hr acutally...

OpenStudy (anonymous):

45 miles / hr is the slower train the express is therefore 60 miles / hr

OpenStudy (anonymous):

oh icic

OpenStudy (anonymous):

i had this exact problem in my homework. you using the holt alg 1 book?

OpenStudy (anonymous):

yhea.... where do u live?

OpenStudy (anonymous):

california

OpenStudy (anonymous):

well durrrr only california uses the california standards

OpenStudy (anonymous):

irvine

OpenStudy (anonymous):

omg me too o.O

OpenStudy (anonymous):

slms?

OpenStudy (anonymous):

lisa?

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

hi :D

OpenStudy (anonymous):

LOL hi

OpenStudy (anonymous):

LOL WHO ARE YOUUUUU

OpenStudy (anonymous):

anthony.... lol, who else did you give this website to?

OpenStudy (anonymous):

AHAHAHAHAHAHHAHA

OpenStudy (anonymous):

HI ANTHONY XDDDDD LOLOLOLOL

OpenStudy (anonymous):

HI LISA xD mr. orchestra?

OpenStudy (anonymous):

no i thought it was you, but ur username is like non anthony ish xD you can tell, mine is lisa ish :D

OpenStudy (anonymous):

yeah... i thought it was you when i read the problem lol i use that username for most websites

OpenStudy (anonymous):

dude i dont get ANY of the math hw TT.TT so as you can see, im posting like everysinge one haha

OpenStudy (anonymous):

o icic xD

OpenStudy (anonymous):

yeah... i noticed i'm not doing my math hw until sunday lol

OpenStudy (anonymous):

well i gtg now cya on monday xD LOL u made my night hahahaha

OpenStudy (anonymous):

hahahaha i wonder what the other person that replied to this post will think XD

OpenStudy (anonymous):

XD